OpenStudy (fibonaccichick666):

For \(p\in Q\) define \(f: R\rightarrow R\) by \(f(x)=[|x|^p sin(\frac{1}{x}) ~~x\not = 0\);\(0~~x=0\). For which values of p is f continuous at 0?

OpenStudy (fibonaccichick666):

@ganeshie8

OpenStudy (fibonaccichick666):

@SolomonZelman , I am not sure if you know this, but could you look at it for me?

OpenStudy (fibonaccichick666):

@freckles do you mind taking a look?

OpenStudy (fibonaccichick666):

I think it is continuous for any value of p>1...

OpenStudy (fibonaccichick666):

but I don't know how to prove it

OpenStudy (freckles):

We want f to be continuous at x=0... we need to have this: \[\lim_{x \rightarrow 0}|x|^psin(\frac{1}{x})=0\] Thinking right now

OpenStudy (fibonaccichick666):

oh, it should include 1. I was thinking differentiable(2nd part of the question)

OpenStudy (freckles):

We know \[\lim_{x \rightarrow 0} x \sin(\frac{1}{x})=0 \text{ by squeeze theroem }\] -1<sin(1/x)<1 -x<x sin(1/x)<x and -x and x go to 0 so x sin(1/x) goes to 0

OpenStudy (fibonaccichick666):

agreed, same for integers above 1

OpenStudy (freckles):

and yeah I didn't put the equal signs there (just the inequalities) because I'm being lazy

OpenStudy (freckles):

Oh true

OpenStudy (freckles):

If you can use squeeze theorem then we are done right? \[-x^n \le \sin(\frac{1}{x} )\le x^{n} \text{ for rational } n \ge 1 \] we have that the thingy in between goes 0 0 since the thingys on the outside go to zero

OpenStudy (fibonaccichick666):

it's the rationals I do not know how to prove. I rewrote p as \(\frac{m}{n}\) then changed \(|x|^\frac{m}{n}=(|x|^m)^\frac{1}{n}\) but I'm not sure that helps

OpenStudy (fibonaccichick666):

can we assume?

OpenStudy (fibonaccichick666):

and yes, we have squeeze up our sleeve

OpenStudy (freckles):

i wonder if we consider rationals between 0 and 1

OpenStudy (fibonaccichick666):

so, jsut cause I'm now thinking about it, what about for 0<p<1

OpenStudy (freckles):

like -x^(1/2) and x^(1/2) also go to zero

OpenStudy (freckles):

as x goes to zero

OpenStudy (fibonaccichick666):

roots are continuous, no?

OpenStudy (fibonaccichick666):

oh but not on R only R+

OpenStudy (fibonaccichick666):

but we have an abs. value so yes, they end up eing continuous

OpenStudy (freckles):

so if we had \[f(x)=x^n \sin(\frac{1}{x}), x \neq 0 \text{ and } 0 ,x =0\] We have by squeeze theorem that when n>0 that \[\lim_{x \rightarrow 0}f(x)=0\] do you agree with this?

OpenStudy (freckles):

The only think I' haven't touched on is the | | part but that shouldn't make much of a difference

OpenStudy (freckles):

i think

OpenStudy (fibonaccichick666):

not in the case of positive exp. No, no difference

OpenStudy (fibonaccichick666):

but yes, I agree

OpenStudy (fibonaccichick666):

|| doesn't effect cont. only differentiability

OpenStudy (freckles):

that is true because |x|->0 as x->0

OpenStudy (freckles):

any |x|^n->0 as x->0 as along as n>0

OpenStudy (freckles):

and not any*

OpenStudy (fibonaccichick666):

? not any? every exponent goes to 0 when x goes to zero.

OpenStudy (fibonaccichick666):

every positive exp*

OpenStudy (freckles):

I was just saying i didn't mean to write any I meant to write and

OpenStudy (fibonaccichick666):

oh ok, you typed any again

OpenStudy (fibonaccichick666):

So now, do you agree that for p>1 it is differentiable?

OpenStudy (fibonaccichick666):

at zero

OpenStudy (freckles):

\[f(x)=x^n \sin(\frac{1}{x}) \text{ if } x>0 , \text{ and } (-x)^n \sin(\frac{1}{x}) \text{ if } x<0,\text{ and } 0 \text{ if } x=0\]

OpenStudy (freckles):

I was going to look at the derivative of this and I think you are right but I have to go right now i have a meeting at 1 oclock sorry

OpenStudy (fibonaccichick666):

it's cool thanks for the help!

OpenStudy (freckles):

we don't care about p<=0 because we already said it wasn't continuous on p<=0 so it certainly won't be differentiable there So we need to look at p>0 only \[f'(x)=nx^{n-1}\sin(\frac{1}{x})+x^n \cdot \frac{-1}{x^2}\cos(\frac{1}{x}) \text{ for } x>0 \\ f'(x)=nx^{n-1}\sin(\frac{1}{x})-x^{n-2}\cos(\frac{1}{x}) \text{ for } x>0 \\ f'(x)=x^{n-2} \cdot [n x \sin(\frac{1}{x})-\cos(\frac{1}{x})] \text{ for } x>0 \\ \text{ and } \\ f'(x)=-n(-x)^{n-1}\sin(\frac{1}{x})+(-x)^{n} \cdot \frac{-1}{x^2}\cos(\frac{1}{x}) \text{ for } x<0 \\ f'(x)=-(-x)^{n-2}[(-x)\sin(\frac{1}{x})+\cos(\frac{1}{x})] \text{ for } x<0\]

OpenStudy (freckles):

we should check to see if the function is actually differentiable at x=0 that is we need to check to see \[\lim_{x \rightarrow 0^+}f'(x)=\lim_{x \rightarrow 0^-}f'(x)=0\] so f'(x)=0 when x=0 so actually n might have to be greater than 2

OpenStudy (freckles):

And I say n might have to be greater than 2 because of x^(n-2) factor